# Interaction-Induced Dimensional Crossover through Full 3D to 1D

### v2

@@ III Equation of state of model system from full 3D to quasi-1D
In the previous Sec. II, we present the novel protocol of realizing interaction-induced dimensional crossover. In Sec. III, we plan to obtain the analytical expressions of equation of state of model system along the dimensional crossover from 3D to quasi-1D.
In this work, we are interested in equation of state at zero temperature. As such, the analytical expression of the ground state energy $E_{g}=\Omega^{0}\left(\mu,n_{0}\right)+\mu N_{0}$ can be obtained as
@@
where the first term represents the mean-field contribution, whereas the second term corresponds to the correction beyond the mean-field approximation, which arises due to quantum fluctuations and the scaling function $f\left(s\right)$ with the variable $s=2J/\tilde{g}_{0}n_{0}$ which can be controlled by the strength of optical lattice, is given by
@@
where the integration over the transverse quasimomenta is restricted to the first Brillouin zone $\left|k_{x}\right|,\left|k_{y}\right|\leq\pi$ [27] and the function ${}_{2}F_{1}\left[a,b,c\text{,}d\right]$ is the hypergeometric function with multipe parameters, including the finite range factor $\chi=4m\tilde{g}_{2}/\hbar^{2}\tilde{g}_{0}$, the dimensionless second-order tunneling rate $\beta=\hbar^{2}J_{1}/2m\tilde{g}_{0}n$ and several combination parameters $\lambda=1+\chi\mu$, $\gamma=1-\cos{k_{x}}/2-\cos{k_{y}}/2$ and $\nu=\left(2+s\gamma+2\chi\mu\cdot\beta\gamma\right)/2\lambda-s\gamma/2$
Quantum depletion refers as to the number of atoms with nonzero momentum which can be calculated through thermodynamic equation $n=-\partial_{\mu}\Omega/V$ [58]. Therefore, we obtain the analytical expression of quantum depletion $\Delta N$ as follows
@@
Equations (11) and (13) are the key results of our work. Based on Eqs. (11) and (13), we investigate the non-universal equations of state of model system along the dimensional crossover from full 3D to quasi-1D with the emphasis on the finite-range interaction effects labelled by $g_{2}$. In this end, we plot the functions of $f\left(s\right)$ and $h\left(s\right)$ of Eqs. (12) and (14) into Figs. 1(a) and (b) respectively, based on which we can analyze interaction-induced dimensional crossover through full 3D to 1D.
@@
In Equation (15), the first term of right-hand side represents the mean-field energy characterized by lattice-renormalized $s$-wave scattering length of $\tilde{a}_{\text{3D}}=a_{\text{3D}}d^{2}/2\pi\sigma^{2}$. The term of $\tilde{a}_{\text{3D}}/\tilde{a}_{\text{cr}}<0$ is a further renormalization of the scattering length due to combined effects of the optical lattice and finite-range interaction. In more details, the $\tilde{a}_{\text{cr}}$ is calculated to be $\tilde{a}_{\text{cr}}=-\frac{d\sqrt{m}}{2\sqrt{2}\left(1.43+9.8\chi\mu\right)\cdot\sqrt{m^{*}}}$. Here, the $m^{*}=\hbar^{2}/Jd^{2}$ is the effective mass associated with the band and the $\chi=4m\tilde{g}_{2}/\hbar^{2}\tilde{g}_{0}$ is due to the finite-range effect. Here, $\tilde{g}_{0}=4\pi\hbar^{2}\tilde{a}_{\text{3D}}/m$ and $\tilde{g}_{2}=2\pi\hbar^{2}\tilde{a}_{\text{3D}}^{2}\tilde{r}_{\text{3D}}/m$ are characterized by the lattice-renormalized $s$-wave scattering length of $\tilde{a}_{\text{3D}}=a_{\text{3D}}d^{2}/2\pi\sigma^{2}$, and the lattice-renormalized effective range of $\tilde{r}_{\text{3D}}=2\pi\sigma^{2}r_{\text{3D}}/d^{2}$. Note that the $\tilde{a}_{\text{cr}}$ can recover the corresponding one in Ref. [27] in the vanishing finite-range effect of $\tilde{g}_{2}=0$.
@@
The last term in Eq. (15) represents the generalized LHY correction in the presence of optical lattice and the finite-rang interaction. We have check that Eq. (15) is simplified into the corresponding result in Ref. [27] for vanishing the finite-range interaction coupling constant $\tilde{g}_{2}=0$; whereas for vanishing the optical lattice $t=0$, our result recovers exactly the corresponding one in Ref. [45]. We see that with respect to the previous cases in Refs. [27, 45], the LHY correction in Eq. (15) is magnified by the renormalization of both the coupling constant of $\tilde{a}_{3D}$, the effective mass of $m^{*}$ and finite-range coupling of $\chi=4m\tilde{g}_{2}/\hbar^{2}\tilde{g}_{0}$.
@@
In the opposite 1D regime corresponding $s\rightarrow 0$, $f\left(s\right)$ exactly approaches a limiting value of $f(0)=4\sqrt{2}/3\left(1+\chi\mu\right)$ with the different values of the finite-range coupling constant of $\chi=4m\tilde{g}_{2}/\hbar^{2}\tilde{g}_{0}$. We routinely check that our calculated $f(0)=4\sqrt{2}/3$ in the case of vanishing $\chi=0$, which can recover the corresponding one in Ref. [27]. In 1D limit, Equation (11) approaches asymptotically to the ground state energy of a 1D Bose gas in the presence of optical lattice and finite-range effect
@@
$\frac{E_{g}}{L}=\frac{1}{2}g_{\text{1D}}n_{\text{1D}}^{2}-\frac{2}{3\pi}\sqrt{m}\left(n_{\text{1D}}g_{\text{1D}}\right)^{\frac{3}{2}}\frac{1}{1+\chi\mu},$ (16)
@@
with $n_{\text{1D}}=n_{0}d^{2}$ being the linear density and $L$ being the length of the tube. In the case of vanishing the finite range interaction coupling constant $\chi=0$, Equation (16) is in agreement with the exact LL solution of the 1D model expanded in the weak coupling regime $mg_{\text{1D}}/\hbar n_{\text{1D}}\ll 1$ [59, 60]. We conclude, therefore, that Eq. (16) is the generalized LL solution of the 1D model expanded in the weak coupling regime in the presence of the finite range interaction.
@@
In a similar fanshion, we delve into the asymptotic behavior of quantum depletion. In the 1D limit of $s\rightarrow 0$, Equqtion (13) diverges as it is expected, indicating that in the absence of tunneling there is no real Bose-Einstein condensation in alignment with the general theorems in one dimension [5]. In the opposite 3D regime of $s\gg 1$, the function of $h\left(s\right)$ in Eq. (14) decays as $\frac{8}{3\sqrt{2}\pi s}\cdot\left(1-3\chi\mu\right)+\frac{4\cdot\left(\chi\mu\right)^{\frac{3}{2}}}{3s}$. Consequently, one finds the analytical expression of the quantum depletion in 3D limit
$n-n_{0}=\frac{8}{3}\sqrt{\frac{\left(\tilde{a}_{\text{3D}}n_{0}\right)^{3}}{\pi}}\frac{m^{*}}{m}\Big{[}1-3\chi\mu+\frac{\sqrt{2}}{2}\pi\left(\chi\mu\right)^{\frac{3}{2}}\Big{]},$ (17)
@@ IV CONCLUSION
In conclusion, we have meticulously derived the equation of state that extends beyond the mean field approximation for a Bose gas with finite-range interactions within a two-dimensional optical lattice. This derivation was achieved using the framework of finite-temperature functional path integrals. Consequently, we have obtained analytical expressions for the ground state energy and quantum depletion. Notably, our work have taken into account finite-range interaction, leading to the dimensional crossover of the system due to constraints on both interaction energy and particle kinetic energy. This crossover differs from the previously observed dimensional crossover in BEC systems, which was solely attributed to the compression of kinetic energy. Our findings provide a more comprehensive understanding of the Bose gas behavior in optical lattices, incorporating the effects of finite-range interactions. Furthermore, in accordance with the Mermin-Wagner-Hohenberg theory [62, 63, 64], quantum fluctuations are significantly enhanced by finite-range corrections in low-dimensional systems. It is due to these reasons that our final results are not universal results which solely depend on $\sqrt{n_{0}a_{s}^{3}}$, as previously believed. Our results demonstrated that the two dimensional optical lattice induces the 3D-1D dimensional crossover in quantum fluctuations. In the 1D limit region, our results closely align with the LL solution, offering a precise representation. Conversely, in the 3D limit region, we obtained the generalized LHY correction. These finite-range analytical results represent nontrivial generalizations of the universal equation of state presented in [27].
The underlying physics of this dimensional crossover involves the intricate interplay among three parameters: the lattice intensity $t$, the effective range $r_{s}$ and the scattering length $a_{s}$. Leveraging the most advanced techniques, all of these quantities are experimentally controllable, allowing for precise manipulation and observation of the system’s behavior. Notably, it is feasible to arbitrarily tune the depth of optical lattices ranging from $0E_{R}$ to $32E_{R}$ [65]. Therefore, given current experimental capabilities, it should be feasible to directly observe phenomena discussed herein, which are intimately linked to dimensional effects. The direct observation of this dimensional effect would mark a significant milestone in elucidating the intricate interplay between dimensionality and quantum fluctuations in low dimensions.

### v3

@@ III Equation of state of model system from full 3D to quasi-1D
In the previous Sec. II, we present the novel protocol for realizing interaction-induced dimensional crossover. In Sec. III, we plan to obtain the analytical expressions of the equation of state of the model system along the dimensional crossover from 3D to quasi-1D.
In this work, we are interested in the equation of state at zero temperature. As such, the analytical expression of the ground state energy $E_{g}=\Omega^{0}\left(\mu,n_{0}\right)+\mu N_{0}$ can be obtained as
@@
where the first term represents the mean-field contribution, whereas the second term corresponds to the correction beyond the mean-field approximation, which arises due to quantum fluctuations and the scaling function $f\left(s\right)$ with the variable $s=2J/\tilde{g}_{0}n_{0}$ which can be controlled by the strength of the optical lattice, is given by
@@
where the integration over the transverse quasimomenta is restricted to the first Brillouin zone $\left|k_{x}\right|,\left|k_{y}\right|\leq\pi$ [28] and the function ${}_{2}F_{1}\left[a,b,c\text{,}d\right]$ is the hypergeometric function with multiple parameters, including the finite range factor $\chi=4m\tilde{g}_{2}/\hbar^{2}\tilde{g}_{0}$, the dimensionless second-order tunneling rate $\beta=\hbar^{2}J_{1}/2m\tilde{g}_{0}n$ and several combination parameters $\lambda=1+\chi\mu$, $\gamma=1-\cos{k_{x}}/2-\cos{k_{y}}/2$ and $\nu=\left(2+s\gamma+2\chi\mu\cdot\beta\gamma\right)/2\lambda-s\gamma/2$
Quantum depletion refers to the number of atoms with nonzero momentum which can be calculated through a thermodynamic equation $n=-\partial_{\mu}\Omega^{\left(0\right)}\left(\mu,n_{0}\right)/V$ with $\Omega^{\left(0\right)}\left(\mu,n_{0}\right)=\Omega_{0}\left(\mu,n_{0}\right)+\Omega_{g}^{\left(0\right)}\left(\mu,n_{0}\right)$. The first term $\Omega_{0}\left(\mu,n_{0}\right)=\left(\frac{1}{2}\tilde{g}_{0}n_{0}^{2}-\mu n_{0}\right)V$ is the mean-field contribution and the second term $\Omega_{g}^{\left(0\right)}\left(\mu,n_{0}\right)=\frac{1}{2}\sum_{\mathbf{k}}E_{\mathbf{k}}\left(\mu,n_{0}\right)$ is the zero temperature contribution of quantum Gaussian fluctuations. From this, we obtain the particle density $n=n_{0}-\frac{1}{2V}\sum_{\mathbf{k}}\partial_{\mu}E_{\mathbf{k}}\left(\mu,n_{0}\right)$. After changing the sum into integrals and using the MCR method again, we obtain the analytical expression of quantum depletion $\Delta N$ as follows
@@
Prior to delving into the intricacies of ground state energy and quantum depletion, it is imperative to examine the rationality of the value of the dimensionless finite-range coupling constant $\chi\mu=4d^{2}/\sigma^{2}\cdot\left(a^{3}_{s}n_{0}\right)\cdot r_{s}/a_{s}$. This step is fundamental as it informs us about the potential experimental viability of our proposed model. Considering the typical experiments in an optically trapped BEC as in [1], the relevant parameter $d/\sigma\sim 1$. To illustrate, let us consider the example of ${}^{6}\text{Li}$, as referenced in [62]. In this instance, the typical density $n_{0}$ is approximately $4\times 10^{12}\text{cm}^{-3}$. Concurrently, the scattering length $a_{s}$ is estimated to be around $1.13\times 10^{-7}\text{m}$. The order of magnitude of $a^{3}_{s}n_{0}$ is approximately $10^{-3}$. Furthermore, as documented in [63], the effective distance $r_{s}$ is estimated to be within the range of $0\sim 3.71\times 10^{-6}\text{m}$. By inserting these parameters into $\chi\mu$, we can approximate its value to be within the range of $0\sim 0.75$. The maximum value of $\chi\mu$ in Fig. 1 is within this range.
Equations (11) and (13) are the key results of our work. Based on Eqs. (11) and (13), we investigate the non-universal equations of state of model system along the dimensional crossover from full 3D to quasi-1D with the emphasis on the finite-range interaction effects labeled by $g_{2}$. In this end, we plot the functions of $f\left(s\right)$ and $h\left(s\right)$ of Eqs. (12) and (14) into Figs. 1(a) and (b) respectively, based on which we can analyze interaction-induced dimensional crossover through full 3D to 1D.
@@
In Equation (15), the first term of the right-hand side represents the mean-field energy characterized by the lattice-renormalized $s$-wave scattering length of $\tilde{a}_{\text{3D}}=a_{\text{3D}}d^{2}/2\pi\sigma^{2}$. The term of $\tilde{a}_{\text{3D}}/\tilde{a}_{\text{cr}}<0$ is a further renormalization of the scattering length due to the combined effects of the optical lattice and finite-range interaction. In more details, the $\tilde{a}_{\text{cr}}$ is calculated to be $\tilde{a}_{\text{cr}}=-\frac{d\sqrt{m}}{2\sqrt{2}\left(1.43+9.8\chi\mu\right)\cdot\sqrt{m^{*}}}$. Here, the $m^{*}=\hbar^{2}/Jd^{2}$ is the effective mass associated with the band and the $\chi=4m\tilde{g}_{2}/\hbar^{2}\tilde{g}_{0}$ is due to the finite-range effect. Here, $\tilde{g}_{0}=4\pi\hbar^{2}\tilde{a}_{\text{3D}}/m$ and $\tilde{g}_{2}=2\pi\hbar^{2}\tilde{a}_{\text{3D}}^{2}\tilde{r}_{\text{3D}}/m$ are characterized by the lattice-renormalized $s$-wave scattering length of $\tilde{a}_{\text{3D}}=a_{\text{3D}}d^{2}/2\pi\sigma^{2}$, and the lattice-renormalized effective range of $\tilde{r}_{\text{3D}}=2\pi\sigma^{2}r_{\text{3D}}/d^{2}$. Note that the $\tilde{a}_{\text{cr}}$ can recover the corresponding one in Ref. [28] in the vanishing finite-range effect of $\tilde{g}_{2}=0$.
@@
The last term in Eq. (15) represents the generalized LHY correction in the presence of the optical lattice and the finite-range interaction. We have checked that Eq. (15) is simplified into the corresponding result in Ref. [28] for vanishing the finite-range interaction coupling constant $\tilde{g}_{2}=0$; whereas for vanishing the optical lattice $t=0$, our result recovers exactly the corresponding one in Ref. [46]. We see that with respect to the previous cases in Refs. [28, 46], the LHY correction in Eq. (15) is magnified by the renormalization of both the coupling constant of $\tilde{a}_{3D}$, the effective mass of $m^{*}$ and the finite-range coupling of $\chi=4m\tilde{g}_{2}/\hbar^{2}\tilde{g}_{0}$.
@@
In the opposite 1D regime corresponding to $s\rightarrow 0$, $f\left(s\right)$ exactly approaches a limiting value of $f(0)=4\sqrt{2}/3\left(1+\chi\mu\right)$ with the different values of the finite-range coupling constant of $\chi=4m\tilde{g}_{2}/\hbar^{2}\tilde{g}_{0}$. We routinely check that our calculated $f(0)=4\sqrt{2}/3$ in the case of vanishing $\chi=0$, which can recover the corresponding one in Ref. [28]. In the 1D limit, Equation (11) approaches asymptotically to the ground state energy of a 1D Bose gas in the presence of the optical lattice and the finite-range effect
@@
$\frac{E_{g}}{L}=\frac{1}{2}g_{\text{1D}}n_{\text{1D}}^{2}-\frac{2}{3\pi}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sqrt{\frac{m}{\hbar^{2}}}}\left(n_{\text{1D}}g_{\text{1D}}\right)^{\frac{3}{2}}\frac{1}{1+\chi\mu},$ (16)
@@
with $n_{\text{1D}}=n_{0}d^{2}$ being the linear density and $L$ being the length of the tube. It is noteworthy to mention that Eq. (16) exhibits a resemblance to the ground state energy of the pure 1D Bose gas in Ref. [40]. However, our study incorporates the presence of an optical lattice and derives the asymptotic energy expression in the 1D limit. The zero-range interaction coupling constant $g_{\text{1D}}$ in Eq. (16) represents the outcome after renormalization. It is indicated that our findings align well with the pure one-dimensional results in the 1D limit. In the case of vanishing the finite range interaction coupling constant $\chi=0$, Equation (16) is in agreement with the exact LL solution of the 1D model expanded in the weak coupling regime $mg_{\text{1D}}/\hbar n_{\text{1D}}\ll 1$ [64, 65]. We conclude, therefore, that Eq. (16) is the generalized LL solution of the 1D model expanded in the weak coupling regime in the presence of the finite range interaction.
@@
In a similar study, we delve into the asymptotic behavior of quantum depletion. In the 1D limit of $s\rightarrow 0$, Equqtion (13) diverges as it is expected, indicating that in the absence of tunneling there is no real Bose-Einstein condensation in alignment with the general theorems in one dimension [5]. In the opposite 3D regime of $s\gg 1$, the function of $h\left(s\right)$ in Eq. (14) decays as $\frac{8}{3\sqrt{2}\pi s}\cdot\left(1-3\chi\mu\right)+\frac{4\cdot\left(\chi\mu\right)^{\frac{3}{2}}}{3s}$. Consequently, one finds the analytical expression of the quantum depletion in 3D limit
$n-n_{0}=\frac{8}{3}\sqrt{\frac{\left(\tilde{a}_{\text{3D}}n_{0}\right)^{3}}{\pi}}\frac{m^{*}}{m}\Big{[}1-3\chi\mu+\frac{\sqrt{2}}{2}\pi\left(\chi\mu\right)^{\frac{3}{2}}\Big{]}.$ (17)
Replace $\chi$ and $\mu$ in Eq. (17) with $4m\tilde{g}_{2}/\hbar^{2}\tilde{g}_{0}$ and $\tilde{g}_{0}n_{0}$, and rewrite Eq. (17) as
$n-n_{0}$ $=$ $n_{0}\Big{[}\frac{8}{3}\frac{m^{*}}{m}\sqrt{\frac{\tilde{a}_{\text{3D}}^{3}n_{0}}{\pi}}-64\frac{m^{*}}{m}\sqrt{\pi}\frac{\tilde{r}_{\text{3D}}}{\tilde{a}_{\text{3D}}}\left(\tilde{a}_{\text{3D}}^{3}n_{0}\right)^{\frac{3}{2}}$ (18)
$+$ $\frac{128}{3}\frac{m^{*}}{m}\pi^{2}\left(\frac{\tilde{r}_{\text{3D}}}{\tilde{a}_{\text{3D}}}\right)^{\frac{3}{2}}\left(\tilde{a}_{\text{3D}}^{3}n_{0}\right)^{2}\Big{]},$
@@ IV CONCLUSION
In conclusion, we have meticulously derived the equation of state that extends beyond the mean field approximation for a Bose gas with finite-range interactions within a two-dimensional optical lattice. This derivation was achieved using the framework of finite-temperature functional path integrals. Consequently, we have obtained analytical expressions for ground state energy and quantum depletion. Notably, our work has taken into account finite-range interaction, leading to the dimensional crossover of the system due to constraints on both interaction energy and particle kinetic energy. This crossover differs from the previously observed dimensional crossover in BEC systems, which was solely attributed to the compression of kinetic energy. Our findings provide a more comprehensive understanding of the Bose gas behavior in optical lattices, incorporating the effects of finite-range interactions. Furthermore, in accordance with the Mermin-Wagner-Hohenberg theory [67, 68, 69], quantum fluctuations are significantly enhanced by finite-range corrections in low-dimensional systems. It is due to these reasons that our final results are not universal results that solely depend on $\sqrt{n_{0}a_{s}^{3}}$, as previously believed. Our results demonstrated that the two dimensional optical lattice induces the 3D-1D dimensional crossover in quantum fluctuations. In the 1D limit region, our results closely align with the LL solution, offering a precise representation. Conversely, in the 3D limit region, we obtained the generalized LHY correction. These finite-range analytical results represent nontrivial generalizations of the universal equation of state presented in [28].
The underlying physics of this dimensional crossover involves the intricate interplay among three parameters: the lattice intensity $t$, the effective range $r_{s}$ and the scattering length $a_{s}$. Leveraging the most advanced techniques, all of these quantities are experimentally controllable, allowing for precise manipulation and observation of the system’s behavior. Notably, it is feasible to arbitrarily tune the depth of optical lattices ranging from $0E_{R}$ to $32E_{R}$ [57]. Therefore, given current experimental capabilities, it should be feasible to directly observe the phenomena discussed herein, which are intimately linked to dimensional effects. The direct observation of this dimensional effect would mark a significant milestone in elucidating the intricate interplay between dimensionality and quantum fluctuations in low dimensions.